OpenStudy (anonymous):

The height of a cylinder is three times the diameter of the base. The surface area of the cylinder is 126pi sq ft. What is the radius of the base?

OpenStudy (tkhunny):

Surface Area of Right Circular Cylinder is ...? In terms of Height and Radius.

OpenStudy (anonymous):

Well. I know the formula is S.A. = L.A.+2B

OpenStudy (anonymous):

Where B is the area of the base.

OpenStudy (tkhunny):

Very good. Now expand both "L.A." and "B" to be in terms of the Radius and Height.

OpenStudy (anonymous):

Would that be 2pi r h + 2pi r squared?

OpenStudy (tkhunny):

Super. One more thing. In the problem statement, we are told h = 3d = 2(3r) = 6r. Do you see this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait I think I get it!

OpenStudy (tkhunny):

Okay, we're ready to solve. Write the rest of the problem statement.

OpenStudy (tkhunny):

See how sly you are?! Now, solve directly for 'r'. Perfect!

OpenStudy (anonymous):

I don't know if I'm doing my math right. I get something not in my choices, but it's closest to 3 ft. So that's what I put. Thank you so much!

OpenStudy (tkhunny):

126pi = 2pi r 6r + 2pi r squared \(2\pi r\cdot 6r + 2\pi r^{2} = 126\pi\) Divide by \(\pi\) \(2 r\cdot 6r + 2r^{2} = 126\) Simplify and divide by 2 \(6r^{2} + r^{2} = 63\) Simplify \(7r^{2} = 63\) Divide by 7 \(r^{2} = 9\) Look's like r = 3 is a good answer. NOT r = -3. That's not a good radius for a real cylinder.

OpenStudy (anonymous):

Did I say -3? But thank you again. I did do my math wrong.

OpenStudy (tkhunny):

No, you did not say "-3". \(r = -3\) IS a solution to \(r^{2} = 9\), but it is not a solution to this problem. Just trying to be thorough.